# Efficiently Vectorized MCMC on Modern Accelerators

### v1

@@ B.1 Algorithms
B.2 FSM Design
Detailed Construction in the two sequential while-loops case
Here, we describe the FSM construction for programs with two sequential while-loops in detail. Following the constructions introduced in the main body, let us note
@@ B.2 FSM Design
Detailed Construction in the two nested while-loop case
Here, we describe the FSM construction for programs with two nested while-loops. Following the constructions introduced in the main body, let us note
@@
    • $\delta_{o}(S_{1},z)=\delta_{o}(S_{i3},z)$ runs the outer while-loop condition on $z$, goes to $S_{i1}$ if True, and $S_{3}$ otherwise.
@@
B.3 Additional Implementation Details
@@
FSM wrapper design. In our experiments, we use a native Python while loop in Algorithm 3, and use a jax.lax.scan to run the FSM step function for blocks of $100$ steps, when drawing $n>100$ samples. This gives us the flexibility to store the results in dynamically shaped lists/arrays and transport to the CPU for faster array slicing when CPU memory is available, whilst still reaping the benefits of JIT compilation.
@@
Compilation. We JIT compile both vmap(step) and vmap(sample) functions for each MCMC algorithm implementation. For Delayed Rejection and the Elliptical Slice Sampler (with $n=25$) we remove compilation time to get more accurate results or the runs with small numbers of chains $m$, due to the low cost of the computations involved.
@@
B.4 Additional Results
@@
Table 2: Effective Sample Size per Second and Kernel Stein Discrepancy (KSD) for transport elliptical slice sampling (TESS (Cabezas and Nemeth,, 2023)) with and without our FSM implementation against NeuTra (Hoffman et al.,, 2019), which also uses preconditioning flows, and two adaptive HMC variants (MEADS (Hoffman and Sountsov,, 2022) and CHEES (Hoffman et al.,, 2021)). Each method uses 128 chains where each chain draws 400 burn-in samples for tuning followed by 1000 warm samples. Our FSM implementation of TESS achieved the best ESS/Second on three out of four problems, in two cases improving the state-of-the-art by roughly a factor of 2.
Dataset Effective Sample Size per Second KSD
MEADS CHEES NeuTra TESS TESS-FSM MEADS CHEES NeuTra TESS TESS-FSM
Predator Prey 1.53 nan 1.59 2.27 3.80 1.13e+6 1e+6 4.5e+5 4.37 7.18
Google Stock 480.76 60.19 185.12 1116.31 2426.16 1.83 0.57 2.73 0.84 0.78
German Credit 141.50 198.99 186.17 58.95 59.04 4.68 6.16 4.58 4.37 4.23
BOD 64.846 247.02 130.59 2978.03 3252.64 2.61e+14 9.63e+15 1.39e+14 49.27 10.80

### v2

@@ B.1 Algorithms
Algorithm 9 bundled_step as input to amortized_step.
1:  Input: Algorithm state $k$, variables $z$
2:  doComputation = False
3:  while not doComputation do
4:     $(k,z,\texttt{isSample},\texttt{doComputation})$$\leftarrow\texttt{step}(k,z)$
5:  end while
6:  Return $(k,z,\texttt{isSample},\texttt{doComputation})$
B.2 FSM Construction Procedure for Programs Considered in Main Text
Detailed Construction in the two sequential while loops case
Here, we describe the FSM construction for programs with two sequential while loops in detail. Following the constructions introduced in the main body, let us note
@@ B.2 FSM Construction Procedure for Programs Considered in Main Text
Detailed Construction in the two nested while loop case
Here, we describe the FSM construction for programs with two nested while loops. Following the constructions introduced in the main body, let us note
@@
    • $\delta_{o}(S_{1},z)=\delta_{o}(S_{i3},z)$ runs the outer while loop condition on $z$, goes to $S_{i1}$ if True, and $S_{3}$ otherwise.
@@
B.3 FSM Construction Procedure for General Programs.
The FSM construction procedures in LABEL:sec:FSM for specific programs can be automated to handle programs with any finite number of sequential and/or nested while loops. We show how this works using a simple programming language defined by the following grammar:
[…]
This grammar contains just enough to describe non-trivial programs containing any finite number of (possibly nested) while loops. Given a program generated with this grammar, we construct an FSM variant of it by:
  1. 1.
    Parsing the program into a parse tree.
  2. 2.
    “Coarsening” the parse tree into a tree where each node is a block.
  3. 3.
    Applying a recursively-defined function which transforms the program into an FSM.
  4. 4.
    Collapsing any spurious empty states.
B.3.1 Step 1: Parsing the program into a parse tree.
First, note that as this grammar is context-free, every program $P$ can be represented by the yield of a parse tree (e.g. Theorem 5.12 (Hopcroft et al.,, 2001)). Consequently, one can use any context-free parser to obtain a parse tree of the program. We choose the look-ahead, left-to-right, rightmost derivation (LALR) parser (DeRemer,, 1969). Figure 8 shows the resulting parse tree associated with the program “op1W{op1W{op1}}”.
Figure 8: Parse tree of the program “op1W{op1W{op1}}”.
B.3.2 Step 2: Coarsening the parse tree.
To convert the parse tree into an FSM, we first derive a simplified version of it with: (i) no operation nodes (e.g. content of simple blocks), (ii) no block markers (e.g. While tokens and braces), (iii) no “blocks” and “block” nodes (these are helpful to clarify the grammar’s definition, but do not carry information about a given program). The first two properties can be achieved by simply pruning the tree of all nodes that are not blocks, which is a well-known operation and not shown here. The output of this procedure is guaranteed to be a tree, as all such nodes (i) and (ii) are terminal. The second property can be achieved by applying the procedure in Algorithm 11.
Algorithm 10 Gathering meaningful children
1:  function gather_children(node) returns list of Node
2:   all_children $\leftarrow$ $\cup$ map(gather_children, node.children)
[…]
4:  if node.type is “block” or “blocks” then
5:      return all_children {Discard the node, and only return the children}
6:  else
7:      return [Node(node.type, all_children)]
8:  end if
9:  end function
Algorithm 11 Coarsening the parse tree
1:  function coarsen_tree(node) returns Node
2:   all_children $\leftarrow$ $\cup$ map(gather_children, children(node))
3:   return Node(“blocks”, all_children)
4:  end function
After applying the coarsening procedure to the parse tree of the same program “op1W{op1W{op1}}”, we obtain the coarsened parse tree shown in Figure 9. The algorithm has been modified to return block labels in order to track the graph manipulation operations done in the next steps.
Figure 9: Coarsened parse tree of the program “op1W{op1W{op1}}”.
B.3.3 Step 3: Transforming the coarsened parse tree into an FSM.
Now we transform the coarsened parse tree into an FSM. Our approach leverages the fact that any subtree of the parse tree is itself a valid (sub)program [^6] It is a “subprogram” of the original program in the sense that it is contained in the original program., and should thus be handle-able by the conversion procedure we define. Each (sub)program is either a single simple block, an arbitrary while block (looping over a sequence of blocks), or the root program (e.g. a sequence of blocks).
This suggests a recursive conversion procedure which computes a single-state FSM for the block at the leaf nodes, and then (recursively) combines these FSMs into larger FSMs at each layer of the parse tree, until we reach the root node, at which point the final FSM is returned. The FSM at leaf nodes (i.e. simple blocks) is a simple single state FSM. The FSM at non-leaf nodes (e.g. while blocks or the entire program) is constructed by (i) connecting the terminal state of the $i^{th}$ child’s FSM to the initial state of the $(i+1)^{th}$ child’s FSM, for every child $i$ of the non-leaf node except the last, and (ii) if the node is a while block, (a) adding an edge from the last child’s FSM to the first child’s FSM, and (b) assigning the initial and terminal state of the current node’s FSM to empty states with an edge between them. Additions (a) and (b) account for the fact that unlike standard blocks, while blocks have bodies that may be looped over, or skipped.
Using empty states as initial and terminal FSM of while blocks allows our FSM conversion procedure to be “context-free”—each subFSM is obtained by assembling sub-subFSMs, without knowledge about the nature of the blocks being connected. Without these empty states, we would require more complicated logic to perform the connection. For instance, in a program of the form “op1W{op1}W{op1}op2”, one needs to connect the first op1 to the last op2 to encode the fact that both while loops may be skipped.
The full algorithm is given in Algorithm 12. The resulting FSM for the program “op1W{op1W{op1}}” is shown in Figure 10. The nodes starting with an “I” (resp. “T”) are the (empty) initial (resp. terminal) states of each while block. The transitions are labeled using the following convention:
[…]
  • “N” means “next block” (which are created when connecting the different subFSMs of the root program)
Figure 10: Intermediate FSM obtained from the coarsened parse tree of the program “op1W{op1W{op1}}”. This intermediate FSM contains empty states, which are collapsed to obtain the final FSM.
Algorithm 12 FSM creation from a coarsened parse tree
1:  function coarse_tree_to_fsm(node) returns FSM
2:  if node.type is “while_block” or “blocks” then
3:      c_fsms $\leftarrow$ map(coarse_tree_to_fsm, node.children)
[…]
7:      // Connect the different subFSMs
[…]
10:      edges, labels $\leftarrow$ edges $\cup$ conn_edges, labels $\cup$ conn_labels
11:      i $\leftarrow$ counter()
12:     if node.type == “while_block” then
13:          terminal_state, initial_state $\leftarrow$ Node(“terminal”), Node(“initial”)
14:          enter_loop_edge $\leftarrow$ (initial_state, c_fsms[0].initial)
15:          exit_loop_edge $\leftarrow$ (c_fsms[-1].terminal, terminal_state)
16:          continue_loop_edge $\leftarrow$ (c_fsms[-1].terminal, c_fsms[0].initial)
17:          skip_loop_edge $\leftarrow$ (terminal_state, initial_state)
[…]
21:          return FSM(nodes, edges, labels)
22:     else
23:          return FSM([simple_block], [ ], [ ])
24:     end if
25:  end if
26:  end function
B.3.4 Step 4: Collapsing the empty states.
The empty states produced by the step above are side effects of the conversion procedure used: the final FSM should not contain them. To this end, we next describe a post-processing, iterative procedure which removes them from the FSM. At each iteration, the procedure selects (if any) an empty state, and—unless it transitions into itself—(i)removes from the list of FSM nodes and (ii) adds edges connecting the parents of the empty state to its children.
The labels of the new edges are obtained by concatenating the “parent-to-empty state” and “empty state-to-children” edges. The procedure is given in Algorithm 13, and the resulting FSM for the program “op1W{op1W{op1}}” is shown in Figure 11. The procedure results in FSMs that agree those in the main text for single while loops, two nested while loops, and two sequential while loops. There are two minor caveats, which we discuss below.
Remark B.1 (Handling empty states with self-transitions).
The procedure above removes all empty states, apart from those with self-transitions. These self-transitions are not present before collapsing the empty states, and arise in the midst of the procedure, after collapsing a subset out of all the FSM empty state. Importantly, such self-transitions are pathological: as the state of the program does not change when performing a self-transition into an empty node, if such a self-transition occurs, it must keep occurring indefinitely afterwards. Thus, programs resulting into FSM with self-recurring empty states form a class of “potentially non-halting” programs.
Remark B.2 (Handling impossible transitions).
The node collapsing process results in the creation of “multistep” edges. These transitions involve evaluating multiple conditions in a single step. Sometimes, these conditions are mutually exclusive: in that case, this multistep transition is impossible, and can be removed from the graph. By removing such transitions after each step of the procedure, the routine can be shown to not add any duplicated edges to the FSM, as proved in the next proposition.
Algorithm 13 Collapsing empty states
1:  function collapse_empty_states(fsm) returns FSM
2:   edges $\leftarrow$ fsm.edges
3:   labels $\leftarrow$ fsm.labels
4:  for all node in filter(empty |$\lvert$| Not initial |$\lvert$| Not final, fsm.nodes) do
5:     if node $\in$ node.children then
6:         continue
7:     end if
8:     for all parent, child in product(node.parents, node.children) do
9:         new_edge $\leftarrow$ (parent, child)
10:        if new_edge $\in$ edges then
11:            error {Edge already exists}
12:        end if
13:         new_label $\leftarrow$ AND(parent.label, child.label)
14:        if Impossible(new_label) then
15:            continue
16:        end if
[…]
19:     end for
20:  end for
21:   nodes $\leftarrow$ filter(non_empty, fsm.nodes)
22:   return FSM(nodes, edges, labels)
23:  end function
Figure 11: FSM for the program “op1W{op1W{op1}}”, with empty states collapsed, using the automatic procedure. Note that this matches the structure of the FSM for the No-U-Turn sampler in Figure 2, which was produced using (essentially) a special case of this procedure for programs with two nested while loops.
B.4 Implementation Details
FSM Step Function Design. At each step of the FSM, one always executes a block $S_{k}:z\mapsto z^{\prime}$ which updates the inputs, before calling the transition function $\delta$ on $(k,S_{k}(z))$ to determine which block should be run next. One can therefore combine these into a single (state) function $T:(k,z)\mapsto(\delta(k,S_{k}(z)),S_{k}(z))$, or equivalently, a collection of functions $T_{1},...,T_{K}$ where $T_{k}:z\mapsto(\delta(k,S_{k}(z)),S_{k}(z))$. In practice, our implementation of Algorithm 2 calls a single switch over these composite state functions to streamline the code. The same composite state functions are also used in our implementation of bundled_step in Algorithm 4.
@@ B.4 Implementation Details
FSM Runtime Design. In our experiments, we use a native Python while loop in Algorithm 3 (and its vmap-ed variant in Algorithm 7). Inside the loop, we use a (JIT compiled) jax.lax.scan to run the FSM step function for blocks of $t=100$ steps, when drawing $n>100$ samples. This gives us the flexibility to store the results in dynamically shaped lists/arrays and transport to the CPU for faster array slicing when CPU memory is available, whilst still reaping the benefits of JIT compilation.
@@
Compilation. We JIT compile both vmap(step) and vmap(sample) functions for each MCMC algorithm implementation (here step refers to any of the basic variant Algorithm 2, bundled variant Algorithm 4, amortized variant Algorithm 5, or the bundled and amortized variant Algorithm 9). For Delayed Rejection and the Elliptical Slice Sampler (with $n=25$) we remove compilation time to get more accurate results or the runs with small numbers of chains $m$, due to the low cost of the computations involved.
@@
Bundling with Amortized Step. As discussed in Section 5, one cannot non-trivially use bundled_step as the step function inside amortized_step, because a given sequence of states may require the amortized computation to be updated in the middle, and nothing in bundled_step flags this. To reap some of the benefits of step bundling when amortizing an expensive computation $g$ (e.g. expensive log-pdf calculations), we use another step function defined in Algorithm 9, as the ‘step’ function called inside amortized_step. This step function iteratively runs step (modified to return the doComputation flag) until doComputtion = True - ironically, using a while loop. When vmap-ed, each ‘step’ inside amortized_step now executes sequences of cheap states that do not require $g$, until $g$ is required again for all chains. This works effectively when $g$ is called inside the iterative state, which is typically the case when $g=\log p$.
B.5 Additional Results
@@
Distribution NUTS-FSM Standard NUTS FSM Speed-up
Real-Estate GPR 863.6 274.1 3.15x
Soil 0.014 0.006 2.43x
Pilots 3.538 3.869 0.91x
Table 2: ESS/sec comparison for NUTS with and without FSM acceleration on Real-Estate GPR sampling problem considered in Section 7.2, and two sampling problems (Soil and Pilots) from the PosteriorDB database (Magnusson et al.,, 2025). Results are averaged over 1000 samples and 128 chains. As with Section 7.4 in the main text, we use 400 warm-up steps to tune the mass matrix and step size of our FSM implementation, and the BlackJAX (non-FSM) implementation. We find substantial speed-ups in 2/3 problems. Both of these problems (Real-estate GPR and Soil) have an expensive log-pdf (and gradient) whilst the log-pdf in Pilots is much cheaper. In general, when the number of integration steps taken by NUTS is moderate to high on average (so that most time is spent integrating along the trajectory), an expensive log-pdf makes the additional cost of executing all other blocks in the FSM step function (when integrating) relatively smaller. That is, the gap between $E(m)$ and $R(m)$ (as defined in Section 4) is smaller, hence increasing the relative efficiency of the FSM. The opposite happens when the log-pdf is cheap (as in Pilots). We also note that computational resources did not balance out across chains in Soil and Pilots after 1000 samples. We therefore would expect to see larger speed-ups when running for longer.